# A stochastic preconditioned Douglas-Rachford splitting method for saddle-point problems

### v2

@@ 2.1 The stochastic preconditioned DR (SPDR) method
With these preparations, now let us turn to the almost sure convergence analysis of (SPDR).
3 Almost sure convergence of stochastic PDR
@@
In this section, we present the almost sure convergence of iteration sequences generated by (SPDR) to a solution of (1.2). To avoid any ambiguity, henceforth, the notation $u^{k}=(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$ along with $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ always represents the stochastic sequences produced by (SPDR). In order to write the iteration in a more compact form, we introduce the notation $\tilde{u}^{k+1}$ for one step of the deterministic update of $u^{k}$, i.e.,
$\tilde{u}^{k+1}:=\mathcal{T}u^{k},\quad\text{with}\ \ \tilde{u}^{k+1}=(\tilde{x}^{k+1},\tilde{y}^{k+1},\tilde{\bar{x}}^{k+1},\tilde{\bar{y}}^{k+1}),$ (3.1)
where $\mathcal{T}$ is the same as (2.4) defined by (PDR). The stochastic update $u^{k+1}$ in (SPDR) can then be written as
$x^{k+1}=\tilde{x}^{k+1},\quad x_{\operatorname{test}}^{k+1}=\tilde{x}_{\operatorname{test}}^{k+1},\quad\bar{x}^{k+1}=\tilde{\bar{x}}^{k+1},$ (3.2)
@@
$y^{k+1}_{i}=\begin{cases}\tilde{y}^{k+1}_{i}&\text{if}\ i=i_{k},\\ y^{k}_{i}&\text{if}\ i\neq i_{k},\end{cases}\quad y^{k+1}_{\operatorname{test},i}=\begin{cases}\tilde{y}^{k+1}_{\operatorname{test},i}&\text{if}\ i=i_{k},\\ y^{k}_{\operatorname{test},i}&\text{if}\ i\neq i_{k},\end{cases}\quad\bar{y}^{k+1}_{i}=\begin{cases}\tilde{\bar{y}}^{k+1}_{i}&\text{if}\ i=i_{k},\\ \bar{y}^{k}_{i}&\text{if}\ i\neq i_{k}.\end{cases}$ (3.3)
Note that $x^{k},x^{k}_{\operatorname{test}},\bar{x}^{k},y^{k},y^{k}_{\operatorname{test}},\bar{y}^{k}$ according to (3.2) and (3.3) are $\mathfrak{F}_{k}$-measurable Borel random vectors by continuity of $\mathcal{T}$ and construction.
The stochastic iteration sequence $\{u^{k}\}$ and the corresponding one-step deterministic update $\tilde{u}^{k+1}$ can be shown in the following diagram
@@
$\downarrow\ \;\;\quad\downarrow\;\;\;\;\quad\downarrow\;\;\;\;\quad\downarrow\;\;\;\qquad\qquad\downarrow{\mathcal{T}}$
$\tilde{u}^{1}\,\ \quad\tilde{u}^{2}\,\ \quad\tilde{u}^{3}\,\ \quad\tilde{u}^{4}\,\ \quad\cdots\,\ \quad\tilde{u}^{k+1}=\mathcal{T}u^{k}\cdots.$
@@
For any $(x,y)\in\operatorname{dom}{\mathcal{F}}\times\operatorname{dom}{\mathcal{G}}$, $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$ and $\bar{y}=y-\tau\mathcal{K}x$, it holds that
$\mathcal{H}\bigl{(}(x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y)\bigr{)}$ $\leq\frac{1}{2}\left(\|u^{k}-u\|^{2}_{\mathbf{M}_{p}}-\mathbb{E}^{k}\|u^{k+1}-u\|^{2}_{\mathbf{M}_{p}}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}\right),$ (3.5)
where $\mathbf{M}_{p}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\tau}(N_{2}-I)\mathbf{P}^{-1},\frac{1}{\sigma}I,\tfrac{1}{\tau}\mathbf{P}^{-1}]$ with $\mathbf{P}=\operatorname{Diag}[p_{1}I,\ldots,p_{n}I]$ and consequently, $\mathbf{P}^{-1}=\operatorname{Diag}[p_{1}^{-1}I,\ldots,p_{n}^{-1}I]$.
@@
With Proposition 1, $u=(x,y,\bar{x},\bar{y})$ and $\tilde{u}^{k+1}$ as in (3.1), (3.2) and (3.3), we arrive at
$\mathcal{H}\bigl{(}(x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y)\bigr{)}$ $\leq\frac{1}{2}\big{(}\|u^{k}-u\|^{2}_{\mathbf{M}}-\|\tilde{u}^{k+1}-u\|^{2}_{\mathbf{M}}-\|\tilde{u}^{k+1}-u^{k}\|^{2}_{\mathbf{M}}\big{)}$ (3.6)
Now, let us reformulate $\|\tilde{y}^{k+1}-\cdot\|_{\frac{(N_{2}-I)}{\tau}}^{2}$ and $\|\tilde{\bar{y}}^{k+1}-\cdot\|_{\tau^{-1}}^{2}$. First, noting that for $\mathfrak{F}_{k}$-measurable Borel random vectors $s^{k}=(s^{k}_{i})_{i=1,\ldots,n}$, $\tilde{s}^{k+1}=(\tilde{s}^{k+1}_{i})_{i=1,\ldots,n}$ on $\mathbb{Y}$, the update rule
@@
Hence, using (3.9) with $\tilde{y}^{k+1}$ and $\tilde{\bar{y}}^{k+1}$ as $\tilde{s}^{k+1}$, $\tfrac{1}{\tau}(N_{2}-I)$ and $\tfrac{1}{\tau}I$ as $\mathbf{B}$, together with the fact that $x^{k+1}$ and $\bar{x}^{k+1}$ are $\mathfrak{F}_{k}$ measurable, we get
$\|\tilde{u}^{k+1}-u\|_{\mathbf{M}}^{2}-\|u^{k}-u\|_{\mathbf{M}}^{2}$ $=\mathbb{E}^{k}\|u^{k+1}-u\|_{\mathbf{M}_{p}}^{2}-\|u^{k}-u\|_{\mathbf{M}_{p}}^{2},$
$\|\tilde{u}^{k+1}-u^{k}\|_{\mathbf{M}}^{2}$ $=\mathbb{E}^{k}\|u^{k+1}-u^{k}\|_{\mathbf{M}_{p}}^{2},$
@@
Let us give some notations first. Denote $\mathcal{Z}^{*}$ as the set for the saddle points of (1.1) and $\mathbb{F}\subset\mathbb{U}$ as the set for the fixed points for $\mathcal{T}$. Introduce $\mathcal{M}_{0}:=\operatorname{Diag}[{N_{1}-I},N_{2}-I,I,I]$ and
$\mathbb{F}_{m}=\mathbb{F}\mathcal{M}_{0}:=\{((N_{1}-I)x^{*},(N_{2}-I)y^{*},\bar{x}^{*},\bar{y}^{*})\ |\ u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})\in\mathbb{F}\}.$
Now, let us turn to the almost sure weak convergence of $\{u^{k}\}_{k}$ to a fixed point of $\mathcal{T}$ as $k\to\infty$. Inspired by [11], we mainly employ the stochastic quasi-Fejér monotonicity [11], Opial’s lemma [2] (Lemma 21), and Lemma 2 to show the almost sure convergence of sequence $\{\tilde{u}^{k+1},k\in\mathbb{N}\}$ first. We then “glue” $\{u^{k+1},k\in\mathbb{N}\}$ to $\{\tilde{u}^{k+1},k\in\mathbb{N}\}$ and finish the proof.
@@
The iteration sequence $\{u^{k}\}_{k}=\{(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})\}_{k}$ produced by (SPDR) weakly converges to a fixed point of $\mathcal{T}$, i.e., $u^{**}=(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})\in\mathbb{F}$ almost surely, and $(x^{**},y^{**})\in\mathcal{Z}^{*}$ is a saddle point for (1.1) with $x^{**}+\sigma\mathcal{K}^{*}y^{**}=\bar{x}^{**}$ and ${y}^{**}-\tau\mathcal{K}{x}^{**}=\bar{y}^{**}$. Moreover, the transitional sequence $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ converges weakly to the saddle point $(x^{**},y^{**})$ almost surely.

### v3

@@ 2 Preconditioned DR framework
The following theorem tells us that over-relaxation can bring certain acceleration for the deterministic PDR method and motivates us to focus on the stochastic and over-relaxed PDR. The proof is given in the Appendix.
Theorem 1.
If $u=(x,y,\bar{x},\bar{y}),$ with $\bar{x}=x+\sigma\mathcal{K}^{\ast}y,$ $\bar{y}=y-\tau\mathcal{K}x$, and $(x,y)$ being a saddle-point of (1.2). Let $u^{k}$ be the iteration sequence generated by RPDR. Denoting $K_{1}=K+1$, let $z^{k}_{\operatorname{test}}=(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ be the transitional variables of RPDR, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K+1}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ converges with rate $\mathcal{O}({1}/{K})$ in a restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, a bounded domain and containing a saddle-point, it holds that
$\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)\leq\frac{1}{2K_{1}\rho_{l}}\sup_{(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{1}}\|u^{0}-u\|_{\mathbf{M}}^{2}=\mathcal{O}(\frac{1}{K}).$ (2.18)
With these preparations for the deterministic relaxed PDR method, i.e., RPDR (or RPDRQ), let us now shift our focus to the stochastic and relaxed PDR method (or stochastic and relaxed PDRQ method).
2.1 The stochastic relaxed preconditioned DR (SRPDR) method
We begin by introducing the stochastic setting. Let $(\Omega,\mathfrak{F},\mathcal{P})$ denote the underlying probability space. Denote $\{i_{0},i_{1},i_{2},\ldots\}$ as index selection variables, i.e., independent random variables on the latter space with values in $\{1,\ldots,n\}$ which are identically distributed. We denote the probability of selecting an index $i\in\{1,\ldots,n\}$ by $p_{i}$ where we assume $10<p_{i}<$ 1 for each $i$ and recall that $\sum_{i=1}^{n}p_{i}=1$. Introduce $P=\operatorname{Diag}[p_{1},p_{2},\ldots,p_{n}]$ and observe that $P^{-1}=\operatorname{Diag}[\frac{1}{p_{1}},\frac{1}{p_{2}},\ldots,\frac{1}{p_{n}}]$. A Borel random vector $y$ in a separable Hilbert space $\mathbb{H}$ is a measurable map from $\Omega$ into $\mathbb{H}$ where $\Omega$ is equipped with $\mathfrak{F}$ and $\mathbb{H}$ is equipped with the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by the open sets of $\mathbb{H}$ (see [15] or [30, Chapter 2]). Further, $(\mathfrak{F}_{k})_{k\in\mathbb{N}}$ denotes the filtration generated by the indices $\{i_{0},i_{1},\ldots,i_{k}\}$, i.e., the $\sigma$-algebras generated by $i_{j}^{-1}\{i\}$ for $j=0,\ldots,k$ and $i=1,\ldots,n$ [40, Chapter 7]. Then, $\mathfrak{F}_{k}\subset\mathfrak{F}_{k+1}$ for each $k\in\mathbb{N}$. Henceforth, we denote $\mathbb{E}^{k}:=\mathbb{E}[\cdot\mid\mathfrak{F}_{k}]$ as the conditional expectation with respect to $\mathfrak{F}_{k}$, and by $\mathbb{E}$ the full expectation. Recall that if a Borel random vector $y$ is $\mathfrak{F}_{k}$-measurable, i.e., the preimages of Borel sets are in $\mathfrak{F}_{k}$, then $\mathbb{E}[y\mid\mathfrak{F}_{k}]=y$.
Let us first extend the relaxed preconditioned Douglas–Rachford splitting method, i.e., RPDR to a stochastic setting. For arbitrary positive step sizes $\sigma$, $\tau$ and $(x^{0},y^{0},\bar{x}^{0},\bar{y}^{0})\in(\mathbb{B}_{1}\times\mathbb{B}_{2})^{2}$, $y^{0}_{\operatorname{test}}\in\mathbb{B}_{2}$, the stochastic relaxed PDR (SRPDR) reads as follows
$\left\{\begin{array}[]{ll}\tilde{x}_{t}^{k+1}=\left(N_{1}+\sigma\tau\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(\left(N_{1}-I\right)x^{k}+\bar{x}^{k}-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right),\\ x^{k+1}=(1-\rho_{k})x^{k}+\rho_{k}\tilde{x}_{t}^{k+1},\\ y^{k+1}=\left\{\begin{array}[]{ll}(1-\rho_{k})y_{i}^{k}+\rho_{k}\tilde{y}^{k+1}_{t,i},\quad\tilde{y}^{k+1}_{t,i}:=N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}\tilde{x}_{t}^{k+1}]&i=i_{k},\\ y^{k}_{i}&\forall i\neq i_{k},\\ \end{array}\right.\\ x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2\tilde{x}_{t}^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}(I+\tau\partial\mathcal{G}_{i})^{-1}[2\tilde{y}^{k+1}_{t,i}-\bar{y}^{k}_{i}]&i=i_{k},\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k},\\ \end{array}\right.\\ \bar{x}^{k+1}=\bar{x}^{k}+\rho_{k,x}(x_{\operatorname{test}}^{k+1}-\tilde{x}_{t}^{k+1}),\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\bar{y}_{i}^{k}+\rho_{k,y}(y_{\operatorname{test},i}^{k+1}-\tilde{y}_{t,i}^{k+1})&i=i_{k},\\ \bar{y}^{k}_{i}&\forall i\neq i_{k},\end{array}\right.\end{array}\right.$ (SRPDR)
where $\tilde{u}^{k+1}_{t}=\mathcal{T}u^{k}:=(\tilde{x}^{k+1}_{t},\tilde{y}^{k+1}_{t},\tilde{\bar{x}}^{k+1}_{t},\tilde{\bar{y}}^{k+1}_{t})$ and $\mathcal{T}$ is the iteration operator of PDR. $N_{1}\succeq I$ and $N_{2}=\operatorname{Diag}[N_{2,1},\ldots,N_{2,n}]$ with each $N_{2,i}:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$ being self-adjoint, bounded and linear such that $N_{2,i}\succeq I$. Again, let $I_{\varrho_{k}}:=\text{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,x}I,\rho_{k,y}I]$, where $\rho_{k}$, $\rho_{k,x}$, and $\rho_{k,y}$ are non-decreasing numbers with lower and upper limits $\rho_{l}$ and $\rho_{u}$, respectively, such that $20<\rho_{l}<\rho_{u}<$ 2. Although the updates of $x^{k+1}$ in the above SRPDR method seem complicated, it is convenient for analysis. Specifically, the updates of $\tilde{x}_{t}^{k+1}$ is based on preconditioned iteration given in (2.2) when $N_{2}=I$, and $N_{1}$ as in (2.1).
To clarify, we refer to the general SRPDR as the stochastic fully relaxed preconditioned Douglas-Rachford method (f-SRPDR) when $I_{\varrho_{k}}=\textrm{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,x}I,\rho_{k,y}I]$, where $\rho_{k}\neq 1$, $\rho_{k,x}\neq 1$, and $\rho_{k,y}\neq 1$. While when $I_{\varrho_{k}}=\textrm{Diag}[I,I,\rho_{k,x}I,\rho_{k,y}I]$, where $\rho_{k,x}\neq 1$, and $\rho_{k,y}\neq 1$, we call it stochastic partial relaxed preconditioned Douglas-Rachford method (p-SRPDR). In addition, by letting $I_{\varrho_{k}}=I$, the stochastic PDR, i.e., SPDR reads as follows
$\left\{\begin{array}[]{ll}x^{k+1}=\left(N_{1}+\sigma\tau\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(\left(N_{1}-I\right)x^{k}+\bar{x}^{k}-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right),\\ y^{k+1}=\left\{\begin{array}[]{ll}N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}x^{k+1}]&i=i_{k},\\ y^{k}_{i}&\forall i\neq i_{k},\\ \end{array}\right.\\ x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}(I+\tau\partial\mathcal{G}_{i})^{-1}[2y^{k+1}_{i}-\bar{y}^{k}_{i}]&i=i_{k},\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k},\\ \end{array}\right.\\ \bar{x}^{k+1}=\bar{x}^{k}+(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}]-x^{k+1},\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\bar{y}_{i}^{k}+y_{\operatorname{test},i}^{k+1}-y_{i}^{k+1}&i=i_{k},\\ \bar{y}^{k}_{i}&\forall i\neq i_{k}.\end{array}\right.\end{array}\right.$ (SPDR)
Considering a simpler case when $N_{1}=I$ and $N_{2}=I$, we will derive the stochastic DR method introduced in [15].
@@ 2.1 The stochastic relaxed preconditioned DR (SRPDR) method
With these preparations, now let us turn to our algorithms’ almost sure convergence and restricted gap estimates. Since we can treat SPDR as a special case of SRPDR, our analysis will focus solely on the convergence properties and gap estimates for SRPDR.
3 Almost sure convergence of stochastic RPDR
@@
This section presents the almost sure convergence of iteration sequences generated by SRPDR to a solution of (1.2). To avoid any ambiguity, henceforth, the notation $u^{k}=(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$ along with $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ always represents the stochastic sequences produced by SRPDR. To write the iteration in a more compact form, we introduce the notation $\tilde{u}^{k+1}$ for one step RPDR update based on $u^{k}$, i.e.,
$\tilde{u}^{k+1}:=\mathcal{T}_{R}u^{k},\quad\text{with}\ \ \tilde{u}^{k+1}=(\tilde{x}^{k+1},\tilde{y}^{k+1},\tilde{\bar{x}}^{k+1},\tilde{\bar{y}}^{k+1}),$ (3.1)
where $\mathcal{T}_{R}$ is the iteration operator of RPDR, and let $(\tilde{x}^{k+1}_{\operatorname{test}},\tilde{y}^{k+1}_{\operatorname{test}})$ be the transitional variables of this one step RPDR update. Further, we introduce the notation $\tilde{u}_{t}^{k+1}$ for one step deterministic PDR update based on $u^{k}$, i.e.,
$\tilde{u}_{t}^{k+1}:=\mathcal{T}u^{k},\quad\text{with}\ \ \tilde{u}_{t}^{k+1}=(\tilde{x}_{t}^{k+1},\tilde{y}_{t}^{k+1},\tilde{\bar{x}}_{t}^{k+1},\tilde{\bar{y}}_{t}^{k+1}),$ (3.2)
where $\mathcal{T}$ is the same as in (2.4), and let $(\tilde{x}^{k+1}_{t,\operatorname{test}},\tilde{y}^{k+1}_{t,\operatorname{test}})$ be the transitional variables of this one step PDR update. Hence the stochastic update $u^{k+1}$ in SRPDR can then be written as
$x^{k+1}=\tilde{x}^{k+1},\quad x_{\operatorname{test}}^{k+1}=\tilde{x}_{\operatorname{test}}^{k+1}=\tilde{x}_{t,\operatorname{test}}^{k+1},\quad\bar{x}^{k+1}=\tilde{\bar{x}}^{k+1},$ (3.3)
@@
$y^{k+1}_{i}=\begin{cases}\tilde{y}^{k+1}_{i}&\text{if}\ i=i_{k},\\ y^{k}_{i}&\text{if}\ i\neq i_{k},\end{cases}\quad y^{k+1}_{\operatorname{test},i}=\begin{cases}\tilde{y}^{k+1}_{\operatorname{test},i}={\tilde{y}^{k+1}_{t,\operatorname{test},i}}&\text{if}\ i=i_{k},\\ y^{k}_{\operatorname{test},i}&\text{if}\ i\neq i_{k},\end{cases}\quad\bar{y}^{k+1}_{i}=\begin{cases}\tilde{\bar{y}}^{k+1}_{i}&\text{if}\ i=i_{k},\\ \bar{y}^{k}_{i}&\text{if}\ i\neq i_{k}.\end{cases}$ (3.4)
Note that $x^{k},x^{k}_{\operatorname{test}},\bar{x}^{k},y^{k},y^{k}_{\operatorname{test}},\bar{y}^{k}$ according to (3.3) and (3.4) are $\mathfrak{F}_{k}$-measurable Borel random vectors by continuity of $\mathcal{T}_{R}$ and construction.
The stochastic iteration sequence $\{u^{k}\}_{k}$ and the corresponding one-step deterministic update $\{\tilde{u}^{k+1}\}_{k+1}$ can be shown in the following diagram
@@
$\downarrow\ \;\;\quad\downarrow\;\;\;\;\quad\downarrow\;\;\;\;\quad\downarrow\;\;\;\qquad\qquad\downarrow{\mathcal{T}_{R}}$
$\tilde{u}^{1}\,\ \quad\tilde{u}^{2}\,\ \quad\tilde{u}^{3}\,\ \quad\tilde{u}^{4}\,\ \quad\cdots\,\ \quad\tilde{u}^{k+1}=\mathcal{T}_{R}u^{k}\cdots.$
@@
For any $u=(x,y,\bar{x},\bar{y})$ with $(x,y)\in\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$ and $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$ and $\bar{y}=y-\tau\mathcal{K}x$, it holds that
$\mathcal{H}\bigl{(}(x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y)\bigr{)}$ $\leq\frac{1}{2}\left(\|u^{k}-u\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\mathbb{E}^{k}\|u^{k+1}-u\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}\right),$ (3.6)
where $\mathbf{M}_{p}:=\operatorname{Diag}[\frac{1}{\sigma}(N_{1}-I),\frac{1}{\tau}(N_{2}-I)\mathbf{P}^{-1},\frac{1}{\sigma}I,\frac{1}{\tau}\mathbf{P}^{-1}]$, and $\mathbf{M}_{p,\varrho_{k}}:=I^{-1}_{\varrho_{k}}\mathbf{M}_{p}$. Here $\mathbf{P}=\operatorname{Diag}[p_{1}I,\ldots,p_{n}I]$, consequently, $\mathbf{P}^{-1}=\operatorname{Diag}[p_{1}^{-1}I,\ldots,p_{n}^{-1}I]$.
@@
First, noting that for $\mathfrak{F}_{k}$-measurable Borel random vectors $s^{k}=(s^{k}_{i})_{i=1,\ldots,n}$, $\tilde{s}^{k+1}=(\tilde{s}^{k+1}_{i})_{i=1,\ldots,n}$ on $\mathbb{Y}$, the update rule
@@
Then with Proposition 2, $u=(x,y,\bar{x},\bar{y})$ and $\tilde{u}^{k+1}$ as in (3.2), (3.3) and (3.4), we arrive at
$\mathcal{H}\bigl{(}(x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y)\bigr{)}$ $\leq\frac{1}{2}\big{(}\|u^{k}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|\tilde{u}^{k+1}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|\tilde{u}^{k+1}-u^{k}\|_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}^{2}\big{)}$ (3.10)
Next, reformulating $\|\tilde{y}^{k+1}-\cdot\|_{(N_{2}-I)\rho^{-1}_{k}\tau^{-1}}^{2}$ and $\|\tilde{\bar{y}}^{k+1}-\cdot\|_{\rho^{-1}_{k,y}\tau^{-1}I}^{2}$ in the above inequality by using (3.9), then together with the fact that $x^{k+1}$ and $\bar{x}^{k+1}$ are $\mathfrak{F}_{k}$ measurable, we get
$\|\tilde{u}^{k+1}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|u^{k}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}$ $=\mathbb{E}^{k}\|u^{k+1}-u\|_{\mathbf{M}_{p,\varrho_{k}}}^{2}-\|u^{k}-u\|_{\mathbf{M}_{p,\varrho_{k}}}^{2},$
$\|\tilde{u}^{k+1}-u^{k}\|_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}^{2}$ $=\mathbb{E}^{k}\|u^{k+1}-u^{k}\|_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}^{2},$
@@
Let us give some notations first. Denote $\mathcal{Z}^{*}$ as the set for the saddle points of (1.1) and $\mathbb{F}\subset\mathbb{U}$ as the set for the fixed points for $\mathcal{T}_{R}$. Now, let us turn to the almost sure weak convergence of $\{u^{k}\}_{k}$ to a fixed point of $\mathcal{T}_{R}$ as $k\to\infty$. Inspired by [15], we mainly employ the stochastic quasi-Fejér monotonicity [15], Opial’s lemma [3, Lemma 21] or [35], and Lemma 2 to show the almost sure convergence of sequence $\{\tilde{u}^{k+1},k\in\mathbb{N}\}$ first. We then “glue” $\{u^{k+1},k\in\mathbb{N}\}$ to $\{\tilde{u}^{k+1},k\in\mathbb{N}\}$ and finish the proof.
@@
The iteration sequence $\{u^{k}\}_{k}=\{(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})\}_{k}$ produced by SRPDR weakly converges to an $\mathbb{F}$-valued random variable $u^{**}=(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ almost surely, and $(x^{**},y^{**})$ is a $\mathcal{Z}^{*}$-valued random variable with $x^{**}+\sigma\mathcal{K}^{*}y^{**}=\bar{x}^{**}$ and ${y}^{**}-\tau\mathcal{K}{x}^{**}=\bar{y}^{**}$. Moreover, the transitional sequence $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ converges weakly to a $\mathcal{Z}^{*}$-valued random variable $(x^{**},y^{**})$ almost surely.